Status Report

NASA ISS On-Orbit Status 23 August 2008

By SpaceRef Editor
August 24, 2008
Filed under , , ,
NASA ISS On-Orbit Status 23 August 2008

All ISS systems continue to function nominally, except those noted previously or below. Saturday – mostly off duty for CDR Volkov, FE-1 Kononenko & FE-2 Chamitoff.

Volkov & Kononenko performed the regular weekly three-hour task of thorough station cleaning. ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the FE’s sleep station with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]

The CDR conducted the periodic service of the active Russian BMP (Harmful Impurities Removal System) by starting the "bake-out" cycle to vacuum on absorbent bed #1 of the regenerable dual-channel filtration system. The process will be terminated at ~5:15pm EDT and Bed #2 regeneration completed tomorrow. [Regeneration of each of the two cartridges takes about 12 hours and is conducted only during crew awake periods. The BMP’s regeneration cycle, normally done every 20 days, is currently performed four times more frequently to remove any lingering Freon-218 from the cabin atmosphere (last time: 8/11&12)];

At ~9:45am EDT the crewmembers conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week’s "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP-Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

For today’s VolSci (Voluntary Weekend Science) program, another session with SHERE (Shear History Extensional Rheology Experiment), FE-2 Chamitoff –

  • Activated the MSG (Microgravity Science Glovebox) from the A31p laptop,
  • Powered on the SHERE hardware,
  • Accessed the CGBA (Commercial Generic Bioprocessing Apparatus) to install the SHERE FM (Fluid Module) #9;
  • Supported the first SHERE experiment run (Test Point 16);
  • Transferred the module with the fluid sample,
  • Installed FM #19 for the second experiment run (Test Point 2);
  • Repeated the procedure (TP 5);
  • Transferred the module with the fluid sample,
  • Installed FM #7 for the third experiment run (Test Point 12);
  • Repeated the procedure (TP 10);
  • Removed the FM from the CGBA, followed by SHERE data transfer;
  • (Had the option of a fourth run if time allowed, with FM #21 (TP 7), plus repeat (TP 13);
  • Turned off the SHERE/CGBA equipment;
  • Transferred the data files to the MSG laptop for subsequent downlink, and
  • Powered down the MSG.

[Background: Rheology is the study of the deformation and flow of matter under the influence of an applied stress (“preshearing” = rotation) which might be, for example, a shear stress or extensional stress. In practice, rheology is principally concerned with extending the "classical" disciplines of elasticity and (Newtonian) fluid mechanics to materials whose mechanical behavior cannot be described with the classical theories. SHERE is designed to study the effect of preshear (rotation) on the transient evolution of the microstructure and viscoelastic tensile stresses for solutions with long chains of monodisperse dilute polymer molecules in the MSG. Collectively referred to as “Boger fluids,” these polymer solutions have become a popular choice for rheological studies of non-Newtonian fluids and are the non-Newtonian fluid used in this experiment. The SHERE hardware consists of the Rheometer, Camera Arm, Interface Box, Cabling, Keyboard, Tool Box, Fluid Modules, and Stowage Tray.]

Kononenko completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]

Gregory filled out the regular FFQ (Food Frequency Questionnaire), his tenth, on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]

The two cosmonauts had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Oleg at ~6:40am, Sergey at ~8:15am EDT.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2), TVIS treadmill (CDR, FE-1), RED resistive exercise device (CDR, FE-1, FE-2) and VELO bike with bungee cord load trainer (CDR, FE-1). Later, Sergey transferred the exercise data file to the MEC laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

As generally every day now, today starting at ~9:00am and running until 3:00pm, the US CDRA (Carbon Dioxide Removal Assembly) was activated intermittently for two half-cycles to control ppCO2 levels. This configuration for the daily ops does not require connecting & disconnecting the ITCS (Internal Thermal Control System) cooling loop. [A forward plan is in work for cycling the CSV (CO2 Selector Valve) to prevent its sticking.]

JEMRMS Update: Yesterday’s Checkout #4 of the JEM Robotic Manipulator System (RMS) was completed for Stage 1 (TVC & Joint Angle Encoder Calibration). Stage 2 (Inching in Manual Mode) had to be aborted due to an emergency stop caused by a time overrun in the ACU (Arm Computer Unit). Investigation by JAXA is underway.

.Weekly Science Update (Expedition Seventeen — Week 18)

3-D SPACE: In progress.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module. Last memory card exchange performed on 8/13.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-2/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): Silicate Garden Hab Side 2 is complete and CGBA-5 has now been powered off awaiting its move to EXPRESS Rack 2 (ER2). The Hab along with other CSI-02 equipment will be removed and returned home on ULF2.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): EuTEF platform is nominal. A DHPU (Data Handling Processing Unit) software patch is currently being developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. On 8/14, EUTEF stayed on feeder#2 (survival heaters) during the duration of the COL S/W Cycle 11 update. The power-up sequence in the evening of 8/14 had some initial problems but could be completed upon a power cycle of the facility. All instruments were subsequently powered up. – DEBIE-2: IOP (Instrument Operations Procedures) performed for 24 hrs duration, after COL S/W Cycle 11 power-down on 8/14-25 and on 8/19-20. Generic status: link error still in work. However a work-around allows for regular science data acquisition using an on-board Instrument Operations Procedure (IOP), but not yet in conjunction with any FIPEX IOP – a final software patch, which will enhance the present instrument capabilities, is under finalization on ground;– DOSTEL: On-going science acquisition;– EuTEMP: Currently inactive as planned;– EVC: Currently no picture taking;– EXPOSE: On-going science acquisition;– FIPEX: IOPs run from 8/15 until 8/18. A new run started on 8/20 and will end on 8/22; — MEDET: Continuous science acquisition; — PLEGPAY: Currently in READY mode, no science data acquisition on-going;– TRIBOLAB: The Ball-Bearing experiment #4 was restarted on 8/18 after thermal stabilization confirmation. This run will be paused for the thruster firing currently planned for 8/27.

FSL (Fluid Science Laboratory): FSL is nominal.

GEOFLOW: The second GEOFLOW run on FSL was performed on 8/19. The temperature gradient set-point of 2.2K could not be reached; actual gradient showed 1.8K. The run, however, was completed nominally with this temperature gradient. Impact on future runs to be assessed by science team. Data downlinking and ground data transfer experienced some difficulties, leading to a delay and cancellation of runs foreseen on 8/19-20. The difficulties are under further investigation, but it is expected that the next GEOFLOW run can be performed on 8/25 upon go-ahead from the science team to erase the data of the first two runs from the FSL hard disk.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Planned.

Marangoni Experiment for ISS (JAXA Fluid Physics Experiment Facility): First Marangoni Experiment onFPEF started on 8/21 during Orbit 2.

Micro-G Clay (JAXA EPO): On 8/11, Greg performed JAXA’s first EPO experiment, Micro-G Clay. The clays are in the drying process.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): Crew performed moon photo activity on 8/19. PI is checking the downlinked images.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-1/-2 (Nitric Oxide Analyzer, ESA): The fourth and last NOA-1 session for both CDR and FE-1 was nominally performed on 8/19.


PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): “Greg, thus far, you have processed 14 Fluid Modules. The minimum number of SHERE test points for this investigation was 9. The preferred number is 20. We also have a significant level of success at 16 Fluid Modules, of which we are short by 2 test points.”

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): In progress.

SOLAR (Solar Monitoring Observatory): Previous Sun observation window closed on 08/08 (GMT221). On 8/14, SOLAR stayed on feeder#2 (survival heaters) during the duration of the COL S/W Cycle 11 update. The power-up sequence in the evening of 8/14 experienced difficulties. Upon a COL LAN switch cycle and SOLAR power cycle, the SOLAR instruments were activated on power feeder#1 again on 8/15. Next sun visibility window to open on 8/25.– SOVIM: out of Sun visibility, so no science data acquisition; — SOLSPEC: out of Sun visibility, so no science data acquisition; — SOLACES: out of Sun visibility, so no science data acquisition.

SOLO: Planned.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.


WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): During the BLB (BIOLAB) checkout after COL S/W (Columbus Orbital Laboratory Software) Cycle 11 update, the Incubator TCS (Thermal Control System) could not be activated. It was suspected that some of the switches on one of the BIOLAB-ESEM (Exchangeable Standard Electronic Modules) boards were out of config. On 8/21, the crew corrected the BLB-ESEM switch positions. This will allow a troubleshooting plan to be developed to recover all the BIOLAB functionalities, following the BIOLAB Rotor A bellow test by ground commanding on 8/22. The next steps in the recovery of all the BIOLAB functionalities will be the replacement of 6 Reference ECs (Experiment Containers) on Rotor B and subsequent test of Rotor B. Planning for these activities is currently TBD.

CEO (Crew Earth Observations): Through 8/21 the ground has received a total of 4,958 frames of CEO images for review and cataloging. “Photos acquired with times corresponding to our CEO daily target list request times are reviewed first and this week included: Mount Rainier (not acquired); Mount Etna (acquired, too cloudy); Mauna Loa, Hawaii (acquired-too cloudy); Andrew Forest (partial acquisition, soft); Santorini Volcanic Complex, Greece (acquired-still under review); Slate Island Impact (acquired); Tunis, Tunisia (acquired-still under review); Libyan megafan (to be reviewed); Algerian megafan (to be reviewed); and Tropical Storm Fay (acquired-captioned and released by PAO). We are seeing improvement in the focus and consistency of your long-lens photos; please continue with more practice when you can. Your rare, striking view of the southern tip of Greenland will be published on NASA/GSFC’s Earth Observatory website this weekend. Your high-oblique, panoramic view was shot as the ISS passed over Newfoundland and offers a great perspective the ice field, glaciers, and topography of part of the Earth’s largest island. Nice shot!”

CEO (Crew Earth Observations) photo targets uplinked for today were Barringer Impact Crater, Arizona (this well-known crater is so young [~50,000 years] that its circular rim has not yet been eroded away. Recent calculations suggest that the impactor was a meteorite ~50m in diameter traveling at 12.8 km/sec [28,600 mph]. Looking right: visual cues are a winding river and the forest line), Jornada Basin, New Mexico (nadir pass. If Greg shot overlapping images of the Rio Grande floodplain [green, vegetated zone] he will have gotten this site. Desertification of local soils, as interpreted from changing vegetation patterns, is the research interest), and Gulf of Fonseca, Central America (looking left for images of this major embayment in the coastline. Three countries meet along the shoreline of the gulf. Heavy use of the coastline includes numerous newly constructed aquaculture basins).

CEO photography can be studied at this “Gateway” website: (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).

Week 19 Scheduled Main Activities:

  • Sun. (8/24): Station cleaning (USOS); BMP ch.2 regen; PFC (FE-2).
  • Mon. (8/25): RPCM LAB1D5/AvRack R&R; O2 port install on WRS; ATV cargo transfers; SVO EDV condensate fill; TEPC relocate; Flowmeter battery maint.; ER5 rack transfer preps.
  • Tue. (8/26): SRVK KAV sample; ATV cargo transfers; ER4/ER5 stow; ER4 relocate to JPM; ER5 relocate to JPM; ER4/ER5 umbillicals mate; ER4/ER5 restow.
  • Wed. (8/27): MBI-12/Sonokard; ATV cargo transfers; LULIN data dwnld.; DC1 IDZ-2 smoke detector cleaning; THC-IMV flow meas.; SVO BRP-M water sampling; IWIS setup; Emergency proc. OBT/drill.

ISS Orbit (as of this morning, 8:50am EDT [= epoch]):
Mean altitude — 355.7 km
Apogee height — 361.6 km
Perigee height — 349.8 km
Period — 91.65 min.
Inclination (to Equator) — 51.64 deg
Eccentricity — 0.000874
Solar Beta Angle — -32.9 deg (magnitude decreasing)
Orbits per 24-hr. day — 15.71
Mean altitude loss in the last 24 hours — 31 m
Revolutions since FGB/Zarya launch (Nov. 98) — 55913

Significant Events Ahead (all dates Eastern Time, some changes possible.):
08/25/08 — Progress M-64/29 SORR (Stage Operations Readiness Review) @ NASA/JSC
09/01/08 — Progress M-64/29P undocking, from FGB nadir; de-orbit/re-entry ~9/8
09/05/08 — ATV1 undocking, from SM aft port (loiter until 9/29 for nighttime reentry/observation)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking (SM aft)
10/01/08 — NASA 50 Years (official)
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft)
10/12/08 — Soyuz TMA-13/17S launch (~3:03am EDT; Lonchakov, Fincke, Garriott)
10/14/08 — Soyuz TMA-13/17S docking (FGB nadir port, ~4:51am)
10/23/08 — Soyuz TMA-12/16S undocking (DC1 nadir) or 10/24?
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/12/08 — STS-126/Endeavour/ULF2 docking
11/20/08 — ISS 10 Years
11/25/08 — Progress M-65/30P undocking & deorbit
11/26/08 — Progress M-66/31P launch
11/30/08 — Progress M-66/31P docking
02/09/09 — Progress M-66/31P undocking & deorbit
02/10/09 — Progress M-67/32P launch
02/12/09 — Progress M-67/32P docking
02/12/09 — STS-119/Discovery/15A launch – S6 truss segment
02/14/09 — STS-119/Discovery/15A docking
02/24/09 — STS-119/Discovery/15A undocking
02/26/09 — STS-119/Discovery/15A landing (nominal)
03/25/09 — Soyuz TMA-14/18S launch
03/27/09 – Soyuz TMA-14/18S docking (DC1)
04/05/09 — Soyuz TMA-13/17S undocking
04/07/09 — Progress M-67/32P undocking & deorbit
05/15/09 — STS-127/Endeavour/2J/A launch – JEM EF, ELM-ES, ICC-VLD
05/25/09 — Soyuz TMA-15/19S launch
05/27/09 — Six-person crew on ISS (following Soyuz 19S docking)
07/30/09 — STS-128/Atlantis/17A – MPLM(P), last crew rotation
10/15/09 — STS-129/Discovery/ULF3 – ELC1, ELC2
12/10/09 — STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 — STS-131/Atlantis/19A – MPLM(P)
04/08/10 — STS-132/Discovery/ULF4 – ICC-VLD, MRM1
05/31/10 — STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

SpaceRef staff editor.